Calculate the Gibbs energy for the next reaction 2NH3 (g) + 2.5O2 (g) = 2NO (g) + 3H2O (g)


1
Expert's answer
2021-01-17T05:46:11-0500




at 298K,

"\Delta G^\circ = \sum {G^\circ _f {\rm{products}}} - \sum {G^\circ _f {\rm{reactants}}}"


ENTHALPY OF REACTION


"{[}4ΔHf(NO (g)) + 6ΔHf(H2O (g)){]} - {[}4ΔHf(NH3
(g)) + 5ΔHf(O2 (g)){]}
\\{[}4(90.25) + 6(-241.82){]} - {[}4(-46.11) +
5(0){]}\\ = -905.48 kJ"



ENTROPY CHANGE 


"{[}4ΔSf(NO (g)) + 6ΔSf(H2O (g)){]} - {[}4ΔSf(NH3 (g)) +
5ΔSf(O2 (g)){]}\\{[}4(210.65) + 6(188.72){]} - {[}4(192.34) + 5(205.03){]}
\\= 180.41 J/K"


FREE ENERGY OF REACTION (AT 298.15 K)

From ΔGf° values:


"{[}4ΔGf(NO (g))
+ 6ΔGf(H2O (g)){]} - {[}4ΔGf(NH3 (g)) + 5ΔGf(O2 (g)){]}\\{[}4(86.57) +
6(-228.59){]} - {[}4(-16.48) + 5(0){]}\\ = -959.34 kJ 
(spontaneous)"


FROM "\Delta G^\circ = \sum {G^\circ _f {\rm{products}}} - \sum {G^\circ _f {\rm{reactants}}}" ,


=-959.27 kJ (spontaneous)